# Zero-shot cross-lingual transfer in instruction tuning of large language models

### v1

@@ Appendix A Extended related work
was extensively studied for discriminative tasks Xue et al. (2021); Conneau et al. (2020); Artetxe et al. (2020); Pires et al. (2019); Wu and Dredze (2019); Pfeiffer et al. (2020) and remains rather under-explored for generative tasks. Vu et al. (2022); Pfeiffer et al. (2023); Maurya et al. (2021); Li and Murray (2023) highlight the problem of generation in the wrong language and propose various approaches to alleviate it. Anonymous (2023) conducts an empirical study of cross-lingual transfer in generation and finds that one of the most important factors enabling transfer is a careful tuning of the learning rate, but focuses on encoder-decoder models and summarization and question answering tasks. In out work we investigate this effect for decoder-only models and in the broader IT setting.
@@
A line of works investigate the native way of achieving instruction following in target languages by using target-language instruction data, obtained by crowd sourcing (Köpf et al., 2023), distillation from strong commercial models (Wei et al., 2023), or automatic translation of English instruction data [^2] https://github.com/avocardio/Zicklein [^3] https://github.com/22-hours/cabrita. Chen et al. (2023) and Kew et al. (2023) focus on compute-efficiency and data-efficiency of multilingual instruction tuning: they highlight the sufficiency of a small amount of updates on multilingual instruction data and of having only three languages in the instruction data, respectively. Ranaldi et al. (2023) propose to include translation-following demonstrations in the instruction data, which are obtained by converting the supervised translation data into the instruction format.
Zhang et al. (2023b) tune the LLM to translate user’s instructions into a pivot language, e.g. English, generate the response in the pivot language and then translate it back into the target language. Such tuning requires access to the instruction data in both target and pivot languages, which is obtained using data translation with ChatGPT.
@@ Appendix B Experimental setup
The main model we study, is LLaMA-2-13B tuned on the Dolly instruction data (15k examples) using full finetuning: LLaMA-2-13B / Dolly-En / FT. LLaMA is a high-quality English-centric model with 2% of pretraining data in languages other than English. This model is release under a License A custom commercial license [^4] https://ai.meta.com/resources/models-and-libraries/llama-downloads/. We also consider several modifications applied to the main model independently one-by-one: reducing model size to 7B (LLaMA-2-7B / Dolly-En / FT), training on a small LIMA data with 1k examples (LLaMA-2-13B / LIMA-En / FT), and adaptation using low-rank adaptation (LoRA) instead of full finetuning ( LLaMA-2-13B / Dolly-En / LoRA).
@@
We create our evaluation set based on AlpacaFarm Dubois et al. (2023), composed of several instruction following test sets. To ensure uniform distribution of tasks in the evaluation set, we identify a diverse set of 25 "tasks" present in AlpacaFarm (e.g. write an email, give home advice, suggest a recipe, etc) and select a subset of 113 instructions from AlpacaFarm that include a balanced number of instructions per "task". For some tasks without enough examples in AlpacaEval, we wrote missing test instructions ourselves. Controlling the task distribution ensures that none of the tasks dominates the evaluation set, leading to more reliable conclusions, and allows us to break down the performance results by tasks, highlighting the types of tasks with high and low performance. A similar strategy of building a balanced over tasks evaluation set was used in Zhang et al. (2023a), but their set is not publicly released.
@@
Figure 4: Per-language results of human evaluation (left columns) and evaluation with GPT-3.5 (right column). All scores from 0 to 2.

### v2

@@ 6 Conclusion
7 Limitations and broader impact
Despite making a substantial effort in systematically evaluating cross-lingual transfer in IT, we acknowledge the infeasibility of considering all possible model configurations and evaluation aspects. First, our study only considers high-resource languages while cross-lingual transfer is expected to pose a greater challenge for medium- and low-resource languages. We focused on high-resource languages as a first step and hope that our evaluation methodology will be helpful in future studies for other language groups. Second, we experiment with one main hyperparameter, learning rate, while other training hyperparameters may also play a substantial role. Nonetheless, we were able to achieve high results even with our rather limited hyperparameter grid. Finally, we only consider commonly used model configurations and adaptations strategies, while other approaches such as reinforcement learning with human feedback, could be also interesting to investigate. We leave their consideration for future work.
We do not anticipate negative societal impact from our work and on the reverse hope that it will help to broaden the accessibility of modern NLP.
@@ Appendix A Extended related work
was extensively studied for discriminative tasks Xue et al. (2021); Conneau et al. (2020); Artetxe et al. (2020); Pires et al. (2019); Wu and Dredze (2019); Pfeiffer et al. (2020) and remains rather under-explored for generative tasks. Vu et al. (2022); Pfeiffer et al. (2023); Maurya et al. (2021); Li and Murray (2023) highlight the problem of generation in the wrong language and propose various approaches to alleviate it. Chirkova and Nikoulina (2024) conduct an empirical study of cross-lingual transfer in generation and finds that one of the most important factors enabling transfer is a careful tuning of the learning rate, but focuses on encoder-decoder models and summarization and question answering tasks. In out work we investigate this effect for decoder-only models and in the broader IT setting.
@@
A line of works investigate the native way of achieving instruction following in target languages by using target-language instruction data, obtained by crowd sourcing (Köpf et al., 2023; Singh et al., 2024), distillation from strong commercial models (Wei et al., 2023; Li et al., 2023), or automatic translation of English instruction data [^2] https://github.com/avocardio/Zicklein [^3] https://github.com/22-hours/cabrita. Chen et al. (2024) and Kew et al. (2023) focus on compute-efficiency and data-efficiency of multilingual instruction tuning: they highlight the sufficiency of a small amount of updates on multilingual instruction data and of having only three languages in the instruction data, respectively. Ranaldi et al. (2023) propose to include translation-following demonstrations in the instruction data, which are obtained by converting the supervised translation data into the instruction format.
Zhang et al. (2023b) tune the LLM to translate user’s instructions into a pivot language, e.g. English, generate the response in the pivot language and then translate it into the target language. Such tuning requires access to the instruction data in both target and pivot languages, which is obtained using data translation with ChatGPT.
@@ Appendix B Experimental setup
The main model we study, is LLaMA-2-13B tuned on the Dolly instruction data (15k examples) using full finetuning: LLaMA-2-13B / Dolly-En / FT. LLaMA is a high-quality English-centric model with 2% of pretraining data in languages other than English. This model is released under a License A custom commercial license [^4] https://ai.meta.com/resources/models-and-libraries/llama-downloads/. We also consider several modifications applied to the main model independently one-by-one: reducing model size to 7B (LLaMA-2-7B / Dolly-En / FT), training on a small LIMA data with 1k examples (LLaMA-2-13B / LIMA-En / FT), and adaptation using low-rank adaptation (LoRA) instead of full finetuning ( LLaMA-2-13B / Dolly-En / LoRA).
@@
We create our evaluation set based on AlpacaFarm Dubois et al. (2023), composed of several instruction following test sets. To ensure uniform distribution of tasks in the evaluation set, we identify a diverse set of 25 "tasks" present in AlpacaFarm (e.g. write an email, give home advice, suggest a recipe, etc) and select a subset of 113 instructions from AlpacaFarm that include a balanced number of instructions per "task". For some tasks without enough examples in AlpacaEval, we wrote missing test instructions ourselves. Controlling the task distribution ensures that none of the tasks dominates the evaluation set, leading to more reliable conclusions, and allows us to break down the performance results by tasks, highlighting the types of tasks with high and low performance. A similar strategy of building a balanced over tasks evaluation set was used in Zhang et al. (2023a).
@@
Figure 4: Per-language results of human evaluation (left columns) and evaluation with GPT-3.5 (right column). All scores from 0 to 2. Heatmap colors visualize written scores.